Thank you for answering this, I appreciate it!!
Problem 9-7A PPE costs; partial years depreciation; alternative methods LO1,2,3 eXcel CHECK FIGURES: 2. $31,320; 3. $21,000 Logic Co. recently negotiated a lump-sum purch of business. The purchase was completed on March 1, 2017, at a total cash price of $1,260,000 a
CHAPTER 9 Property. Plant and Equipment and Intangibles The estimated market valu 100, land improvements, $68,000, and vehicles, $176.8 included a building, land, certain land improvements, and 12 vehicles. the assets were building, $652,800; land, $462,400; land improvements The companys fiscal year ends on December 31. Required 1. Prepare a schedule to allocate the lump-sum purchase price to the sepa $462,4 chased. Also present the journal entry to record the purchase. 2. Calculate the 2017 depreciation expense on the building using the straig est whole month, assuming a 15-year life and a $41,040 residual value culate the 2017 depreciation expense on the land improvements assuming a five-year life, residual, and double-declining-balance depreciation calculated to the nearest whole 3. Cal month. Analysis Component: Assum 23, 2017. Would this affect your answers in parts 2 and 3 above? Explain. e the assets purchased on March 1, 2017, were not put into service until May Req
Answer #1
1 Assets Market
value
% of total
market
value
* Total
cost of
acquisition
= Allocated
purchase
price
Building 652800 48% 1260000 604800
Land 462400 34% 1260000 428400
Land improvements 68000 5% 1260000 63000
Vehicles 176800 13% 1260000 163800
1360000 1260000
Journal entry:
Date General journal Debit Credit
Mar 1. Building 604800
Land 428400
Land improvements 63000
Vehicles 163800
Cash 1260000
2 Depreciation expense using straight line method=(Cost - Residual value)/Useful life=(604800-41040)/15=$ 37584
But, the building purchased on March 1. Hence,depreciation to be computed for Mach to December (10 months)
Depreciation for 2017=37584*(10/12)=$ 31320
3 Depreciation rate as per double declinig balance method=2*(1/useful life)*100=2*(1/5)*100=40%
Depreciation for 2017-10 months (mar to dec)=cost*Depreciation rate=63000*40%*(10/12)=$ 21000
Aalysis:
Depreciaton to be computed from the when asset is ready for use and not when the asset was actually put to use
Here, Asset purchased on march 1 and and put to service on May 23.
Hence,depreciation to be computed from march 1.
This will not affect answers in part 2 and 3.